# The Price of Competitive Information Disclosure

### v1

@@ 5.2 Case 1: All Agents in $\mathcal{N}_{2}$ are Large Contributors
$\mathrm{SW}(\bm{\mathcal{Z}})\geq\mathrm{SW}^{\prime}\cdot\frac{\beta\cdot(Q-1/\alpha)}{1+Q}.$
@@
$\mathrm{SW}(\bm{\mathcal{Z}})$ $=\sum_{i\in[N]}c_{i}\geq\sum_{i\in\mathcal{N}_{2}}c_{i}\geq\beta\cdot\sum_{i\in\mathcal{N}_{2}}E_{i}\cdot q_{i}$
@@
$\geq\beta\cdot\mathrm{SW}^{\prime}\cdot\frac{Q}{1+Q}\cdot\frac{Q-1/\alpha}{Q}=\mathrm{SW}^{\prime}\cdot\frac{\beta\cdot(Q-1/\alpha)}{1+Q}.$
@@
Since $\mathrm{SW}^{\max}\leq\mathrm{SW}^{\prime}$, it follows that the Price of Anarchy is bounded by $\frac{1+Q}{\beta\cdot(Q-1/\alpha)}$ in this case.
@@ 5.3 Case 2: Some Agent in $\mathcal{N}_{2}$ Is a Small Contributor
$c_{i^{*}}<E\cdot\beta\cdot q_{i^{*}}.$
@@
[…]
@@ 5.3.1 Construction of the deviation signal $s^{*}$
$\pi_{i^{*}}^{\prime}(S\mid v)=\begin{cases}\pi_{i^{*}}(S\cap S_{1}\mid v)+\pi_{i^{*}}(S_{2}\mid v)\cdot\mathbb{1}[s^{*}\in S]&\text{if $v$ is in the top $\phi_{i^{*}}$-quantile of $\mathcal{F}_{i^{*}}$,}\\ \pi_{i^{*}}(S\mid v)&\text{otherwise.}\end{cases}$
@@
Figure 2: Illustration of construction of deviation signaling scheme $s^{*}$.
Fig. 2 illustrates the process of the construction of $s^{*}$. The $x$-axis represents the values while the $y$-axis (heights of the bars) represents the mass on the corresponding values. Since each signal consists of masses on different values, we use the same color set of rectangles to represent a signal. In the illustration, we assume $\Sigma_{i^{*}}$ consists of five signals $\{s_{i^{*}}^{j}\}_{j=1}^{5}$. Assume that if signal $s_{i^{*}}^{1}$ or $s_{i^{*}}^{4}$ is sent, agent $i^{*}$ has at least $\tau$ probability of being selected. Then $s_{i^{*}}^{1},s_{i^{*}}^{4}\in S_{1}$. On the other hand, $s_{i^{*}}^{2}$, $s_{i^{*}}^{3}$, or $s_{i^{*}}^{5}$ belong to $S_{2}$. We will pick all the masses from by these signals with values in the top $\phi_{i^{*}}$ quantile, and combine them into a signal $s^{*}$, as shown in the tiled area on the right.
@@ 5.3.2 The posterior mean of $s^{*}$
Since $c_{i^{*}}<E_{i^{*}}\cdot\beta\cdot q_{i^{*}}$, the value contribution comes from signals in $S_{1}$ or signals in $S_{2}$. Therefore, by considering only signals in $S_{1}$, we have
@@
By definition of $S_{1}$, for all $\sigma\in S_{1}$, $w_{i^{*}}(\sigma)>\tau$. We have
@@
Since $s^{*}$ is sent only when agent $i^{*}$’s value is in the top $\phi_{i^{*}}$ quantile, it must hold that 0<Pr[s is $0<\Pr[\textrm{$s^{*}$ is sent}]\leq\phi\cdot q_{i^{*}}$. Combining with Eqs. 4 and 5, we have
@@ 5.3.3 Deviation to $s^{*}$ in new signaling scheme
    If $\phi_{i^{*}}=\phi\cdot q_{i^{*}}$, consider the case after agent $i^{*}$ deviates to $\mathcal{Z}_{i^{*}}^{\prime}$. All values in the top $(\phi\cdot q_{i^{*}})$ quantile are mapped to either signals in $S_{1}$ or $s^{*}$. Since when any signal in $S_{1}$ or $s^{*}$ is sent to the principal, agent $i^{*}$ is selected with probability at least $\tau$, the deviation to $s^{*}$ makes her overall probability of being selected at least $\phi\cdot q_{i^{*}}\cdot\tau$. By the construction of $\mathcal{Z}_{i^{*}}$, agent $i^{*}$ loses utility from the cases when her value is at most $v_{i^{*}}(\phi_{i^{*}})$ and gains utility from the cases when her value is at least $v_{i^{*}}(\phi_{i^{*}})$. Recall that $i^{*}\in\mathcal{N}_{2}$, so $r_{i^{*}}(\bm{\mathcal{Z}})<\alpha\cdot q_{i^{*}}$. Since the parameters satisfy $\phi\cdot\tau>\alpha$, agent $i^{*}$’s overall probability of being selected increases from strictly less than $\alpha\cdot q_{i^{*}}$ to $\phi\cdot\tau\cdot q_{i^{*}}$. She will gain at least
    $(\phi\cdot\tau-\alpha)\cdot q_{i^{*}}\cdot u_{i^{*}}(v_{i^{*}}(\phi_{i^{*}}))>0$
@@
Since $E_{i^{*}}\geq E_{\mathrm{cut}}$, all agents in $\mathcal{N}_{\ell}$ must have their $q_{i}=1$. Note that $i^{*}\in\mathcal{N}_{s}$. Since $q_{i^{*}}>0$ and $\sum_{i\in[N]}q_{i}=Q\cdot k$, we have $|\mathcal{N}_{\ell}|<k\cdot Q\leq k$, where we have used $Q\leq 1$.
@@
If not, we have $\Pr[M(\bm{\mathcal{Z}})\geq e_{i^{*}}(s^{*})]<1-\tau\Rightarrow\Pr[M(\bm{\mathcal{Z}})<e_{i^{*}}(s^{*})]\geq\tau$. Denote by $M_{-i^{*}}(\bm{\mathcal{Z}})$ the random variable representing the $(k-|\mathcal{N}_{\ell}|)$-th largest posterior mean among all agents except $i^{*}$ in $\mathcal{N}_{s}$ in $\bm{\mathcal{Z}}$. $M_{-i^{*}}(\bm{\mathcal{Z}})$ is stochastically dominated by $M(\bm{\mathcal{Z}})$. When $M_{-i^{*}}(\bm{\mathcal{Z}})<e_{i^{*}}(s^{*})$, agent $i^{*}$ must be selected when $s^{*}$ is sent. This contradicts the assumption that she is selected with probability less than $\tau$ when $s^{*}$ is selected. ∎
@@
Given a strategy profile $\bm{\mathcal{Z}}$, if $\Pr\left[M(\bm{\mathcal{Z}})\geq e_{i^{*}}(s^{*})\right]\geq 1-\tau$, then we have
$\mathrm{SW}(\bm{\mathcal{Z}})\geq\mathrm{SW}^{\prime}(\bm{\mathcal{Z}})\cdot\frac{(1-\tau)\cdot\left(\frac{1}{\phi}-\frac{\beta}{\tau\cdot\phi}\right)}{1+Q}.$
@@
$\mathrm{SW}^{\textrm{(fake)}}(\bm{\mathcal{Z}})$ $\geq(k-|\mathcal{N}_{l}|)\cdot\mathbb{E}[M(\bm{\mathcal{Z}})]+\sum_{i\in\mathcal{N}_{\ell}}E_{i}$
$\geq(1-\tau)\cdot(k-|\mathcal{N}_{l}|)\cdot e_{i^{*}}(s^{*})+\sum_{i\in\mathcal{N}_{\ell}}E_{i}$ (w.p. $1-\tau$, $M(\bm{\mathcal{Z}})\geq e_{i^{*}}(s^{*})$. )
@@
To see the last inequality, note that $Q\leq 1$, so that $\alpha>\frac{1}{Q}\geq 1$. Since $\tau<1$, we have $\phi\geq\frac{\alpha}{\tau}>1$. Since $\beta\leq\tau$, we finally have
@@
By letting $b=\frac{\sum_{i\in\mathcal{N}_{\ell}}E_{i}}{|\mathcal{N}_{\ell}|}$, $w_{b}=|\mathcal{N}_{\ell}|$, $a_{\ell}=E_{i^{*}}$, $w_{s}=(1-\tau)\cdot\left(\frac{1}{\phi}-\frac{\beta}{\tau\cdot\phi}\right)\cdot(k-|\mathcal{N}_{\ell}|)$, $a_{s}=\frac{k\cdot E_{\mathrm{cut}}+\sum_{i\in\mathcal{N}_{s}}E_{i}\cdot q_{i}}{k+\sum_{i\in\mathcal{N}_{s}}q_{i}}$, $w_{\ell}=k+\sum_{i\in\mathcal{N}_{s}}q_{i}$, it follows from Eq. 6 that
@@
$=\frac{\mathrm{SW}^{\prime}}{(1+Q)\cdot k}\cdot(1-\tau)\cdot\left(\frac{1}{\phi}-\frac{\beta}{\tau\cdot\phi}\right)\cdot k.$
@@
Since $\mathrm{SW}^{\prime}$ is an upper bound on social welfare in the full-revelation scenario, and since selection in the hypothetical scenario leads to at most the social welfare of selecting the $k$ agents with highest posterior mean, the PoA in this case is at most $\frac{1+Q}{(1-\tau)\cdot\left(\frac{1}{\phi}-\frac{\beta}{\tau\cdot\phi}\right)}$.
Completing the proof of Theorem 5.1.
@@
$\mbox{PoA}\leq\max\left\{\frac{1+Q}{\beta\cdot(Q-1/\alpha)},\frac{1+Q}{(1-\tau)\cdot\left(\frac{1}{\phi}-\frac{\beta}{\tau\cdot\phi}\right)}\right\}.$
By setting
$Q=1,\qquad\alpha=\frac{\sqrt{5}+1}{2}\approx 1.618,\qquad\beta=\sqrt{5}-2\approx 0.236,$
@@
we know that the PoA of any instance is at most $11+5\sqrt{5}\approx 22.18$. This completes the proof of Theorem 5.1. ∎

### v2

@@ 5.2 Case 1: All Agents in $\mathcal{N}_{2}$ are Large Contributors
$\mathrm{SW}(\bm{\mathcal{Z}})\geq\mathrm{SW}^{\prime}\cdot\frac{\beta\cdot(1-1/\alpha)}{2}.$
@@
$\mathrm{SW}(\bm{\mathcal{Z}})$ $=\sum_{i\in[N]}c_{i}(\bm{\mathcal{Z}})\geq\sum_{i\in\mathcal{N}_{2}}c_{i}(\bm{\mathcal{Z}})\geq\beta\cdot\sum_{i\in\mathcal{N}_{2}}E_{i}\cdot q_{i}$
@@
$\geq\beta\cdot\mathrm{SW}^{\prime}\cdot\frac{k}{1+1}\cdot\frac{1-1/\alpha}{k}=\mathrm{SW}^{\prime}\cdot\frac{\beta\cdot(1-1/\alpha)}{2}.$
@@
Since $\mathrm{SW}^{\max}\leq\mathrm{SW}^{\prime}$, it follows that the Price of Anarchy is bounded by $\frac{2}{\beta\cdot(1-1/\alpha)}$ in this case.
@@ 5.3 Case 2: Some Agent in $\mathcal{N}_{2}$ Is a Small Contributor
$c_{i^{*}}<E_{i^{*}}\cdot\beta\cdot q_{i^{*}}.$
@@
Fix some constant $\phi>1$ and consider the top $\phi_{i^{*}}=\min\{\phi\cdot q_{i^{*}},\ 1\}$ quantile of $i^{*}$’s value distribution. Let $\bm{\Pi}_{-i^{*}}=\bigotimes_{j\neq i^{*}}\Pi_{j}$ denote the distribution of signals from all agents except $i^{*}$. For simplicity, for any signal profile $\bm{\sigma}$, we write $\rho_{i^{*}}(\bm{\sigma})$ as $\rho_{i^{*}}(\sigma_{i^{*}},\bm{\sigma}_{-i^{*}})$ where $\bm{\sigma}_{-i^{*}}$ is the signal profile of all agents except $i^{*}$. For each signal $\sigma_{i^{*}}\in\Sigma_{i^{*}}$, define
@@ 5.3.1 Construction of the deviation signal $s^{*}$
[…]
@@
Figure 2: Illustration of construction of deviation signaling scheme with signal $s^{*}$.
Fig.˜2 illustrates the process of the construction of $s^{*}$. The $x$-axis represents the values while the $y$-axis (heights of the bars) represents the mass on the corresponding values. Since each signal consists of masses on different values, we use the same color set of rectangles to represent a signal. In the illustration, we assume $\Sigma_{i^{*}}$ consists of five signals $\{s_{i^{*}}^{j}\}_{j=1}^{5}$. Assume that if signal $s_{i^{*}}^{1}$ or $s_{i^{*}}^{4}$ is sent, agent $i^{*}$ has at least $\tau$ probability of being selected. Then $s_{i^{*}}^{1},s_{i^{*}}^{4}\in S_{1}$. On the other hand, $s_{i^{*}}^{2}$, $s_{i^{*}}^{3}$, or $s_{i^{*}}^{5}$ belong to $S_{2}$. We will pick all the masses from these signals with values in the top $\phi_{i^{*}}$ quantile, and combine them into a signal $s^{*}$, as shown in the tiled area on the right.
@@ 5.3.2 The posterior mean of $s^{*}$
Since $c_{i^{*}}<E_{i^{*}}\cdot\beta\cdot q_{i^{*}}$ (by the initial assumption of this case) and the value contribution comes from signals in $S_{1}$ or signals in $S_{2}$, by considering only signals in $S_{1}$, we have
@@
By definition of $S_{1}$, for all $\sigma\in S_{1}$, $w_{i^{*}}(\sigma)\geq\tau$. We have
@@
Since $s^{*}$ is sent only when agent $i^{*}$’s value is in the top $\phi_{i^{*}}$ quantile, it must hold that $0<\Pr[\textrm{$s^{*}$ is sent}]\leq\phi\cdot q_{i^{*}}$. Combining with Eqs.˜4 and 5, we have
@@ 5.3.3 Deviation to $s^{*}$ in new signaling scheme
    If $\phi_{i^{*}}=\phi\cdot q_{i^{*}}$, consider the case after agent $i^{*}$ deviates to $\mathcal{Z}_{i^{*}}^{\prime}$. All values in the top $(\phi\cdot q_{i^{*}})$ quantile are mapped to either signals in $S_{1}$ or $s^{*}$. Since when any signal in $S_{1}$ or $s^{*}$ is sent to the principal, agent $i^{*}$ is selected with probability at least $\tau$, the deviation to $s^{*}$ makes her overall probability of being selected at least $\phi\cdot q_{i^{*}}\cdot\tau$. By the construction of $\mathcal{Z}_{i^{*}}^{\prime}$, agent $i^{*}$ loses utility from the cases when her value is at most $v_{i^{*}}(1-\phi_{i^{*}})$ and gains utility from the cases when her value is at least $v_{i^{*}}(1-\phi_{i^{*}})$. Recall that $i^{*}\in\mathcal{N}_{2}$, so $r_{i^{*}}(\bm{\mathcal{Z}})<\alpha\cdot q_{i^{*}}$. Since the parameters satisfy $\phi\cdot\tau>\alpha$, agent $i^{*}$’s overall probability of being selected increases from strictly less than $\alpha\cdot q_{i^{*}}$ to $\phi\cdot\tau\cdot q_{i^{*}}$. She will gain at least
    $(\phi\cdot\tau-\alpha)\cdot q_{i^{*}}\cdot u_{i^{*}}(v_{i^{*}}(1-\phi_{i^{*}}))>0$
@@
Since $E_{i^{*}}\geq E_{\mathrm{cut}}$, all agents in $\mathcal{N}_{\ell}$ must have their $q_{i}=1$. Note that $i^{*}\in\mathcal{N}_{s}$. Since $q_{i^{*}}>0$ and $\sum_{i\in[N]}q_{i}=k$, we have $|\mathcal{N}_{\ell}|<k$.
@@
If not, we have $\Pr[M(\bm{\mathcal{Z}})\geq e^{\prime}_{i^{*}}(s^{*})]<1-\tau$ and thus $\Pr[M(\bm{\mathcal{Z}})<e^{\prime}_{i^{*}}(s^{*})]\geq\tau$. Denote by $M_{-i^{*}}(\bm{\mathcal{Z}})$ the random variable representing the $(k-|\mathcal{N}_{\ell}|)$-th largest posterior mean among all agents except $i^{*}$ in $\mathcal{N}_{s}$ in $\bm{\mathcal{Z}}$. $M_{-i^{*}}(\bm{\mathcal{Z}})$ is stochastically dominated by $M(\bm{\mathcal{Z}})$. When $M_{-i^{*}}(\bm{\mathcal{Z}})<e^{\prime}_{i^{*}}(s^{*})$, agent $i^{*}$ must be selected when $s^{*}$ is sent. This contradicts the assumption that she is selected with probability less than $\tau$ when $s^{*}$ is selected. ∎
@@
Given a strategy profile $\bm{\mathcal{Z}}$, if $\Pr\left[M(\bm{\mathcal{Z}})\geq e^{\prime}_{i^{*}}(s^{*})\right]\geq 1-\tau$, then we have
$\mathrm{SW}(\bm{\mathcal{Z}})\geq\mathrm{SW}^{\prime}\cdot\frac{(1-\tau)\cdot\left(\frac{1}{\phi}-\frac{\beta}{\tau\cdot\phi}\right)}{2}.$
@@
$\mathrm{SW}^{\textrm{(fake)}}(\bm{\mathcal{Z}})$ $\geq(k-|\mathcal{N}_{\ell}|)\cdot\mathbb{E}[M(\bm{\mathcal{Z}})]+\sum_{i\in\mathcal{N}_{\ell}}E_{i}$
$\geq(1-\tau)\cdot(k-|\mathcal{N}_{\ell}|)\cdot e^{\prime}_{i^{*}}(s^{*})+\sum_{i\in\mathcal{N}_{\ell}}E_{i}$ (w.p. $1-\tau$, $M(\bm{\mathcal{Z}})\geq e^{\prime}_{i^{*}}(s^{*})$. )
@@
To see the last inequality, note that $\alpha>1$. Since $\tau<1$, we have $\phi\geq\frac{\alpha}{\tau}>1$. Since $\beta\leq\tau$, we finally have
@@
By letting $b=\frac{\sum_{i\in\mathcal{N}_{\ell}}E_{i}}{|\mathcal{N}_{\ell}|}$, $w_{b}=|\mathcal{N}_{\ell}|$, $a_{\ell}=E_{i^{*}}$, $w_{s}=(1-\tau)\cdot\left(\frac{1}{\phi}-\frac{\beta}{\tau\cdot\phi}\right)\cdot(k-|\mathcal{N}_{\ell}|)$, $a_{s}=\frac{k\cdot E_{\mathrm{cut}}+\sum_{i\in\mathcal{N}_{s}}E_{i}\cdot q_{i}}{k+\sum_{i\in\mathcal{N}_{s}}q_{i}}$, $w_{\ell}=k+\sum_{i\in\mathcal{N}_{s}}q_{i}$, it follows from Eq.˜6 that
@@
$=\frac{\mathrm{SW}^{\prime}}{2k}\cdot(1-\tau)\cdot\left(\frac{1}{\phi}-\frac{\beta}{\tau\cdot\phi}\right)\cdot k.$ (Since $q_{i}=1$ for any $i\in\mathcal{N}_{\ell}$)
@@
Since $\mathrm{SW}^{\prime}$ is an upper bound on social welfare in the full-revelation scenario, and since selection in the hypothetical scenario leads to at most the social welfare of selecting the $k$ agents with highest posterior mean, the PoA in this case is at most $\frac{2}{(1-\tau)\cdot\left(\frac{1}{\phi}-\frac{\beta}{\tau\cdot\phi}\right)}$.
Completing the proof of Theorem˜5.1.
@@
$\mbox{PoA}\leq\max\left\{\frac{2}{\beta\cdot(1-1/\alpha)},\frac{2}{(1-\tau)\cdot\left(\frac{1}{\phi}-\frac{\beta}{\tau\cdot\phi}\right)}\right\}.$
By setting [^16] The current parameters are set with $\phi\cdot\tau=\alpha$, we can jitter the parameters a bit to achieve $\phi\cdot\tau>\alpha$ as required. So the PoA bound can be arbitrarily close to $11+5\sqrt{5}$.
$\alpha=\frac{\sqrt{5}+1}{2}\approx 1.618,\qquad\beta=\sqrt{5}-2\approx 0.236,$
@@
we know that the PoA of any instance is at most $11+5\sqrt{5}\approx 22.181$. This completes the proof of Theorem˜5.1. ∎